# Region-Wise Attentive Multi-View Representation Learning for Urban Region Embeddings

### v1

@@ Abstract.
Urban region embedding is an important and yet highly challenging issue due to the complexity and constantly changing nature of urban data. To address the challenges, we propose a Region-Wise Multi-View Representation Learning (ROMER) to capture multi-view dependencies and learn expressive representations of urban regions without the constraints of rigid neighbourhood region conditions. Our model focus on learn urban region representation from multi-source urban data. First, we capture the multi-view correlations from mobility flow patterns, POI semantics and check-in dynamics. Then, we adopt global graph attention networks to learn similarity of any two vertices in graphs. To comprehensively consider and share features of multiple views, a two-stage fusion module is further proposed to learn weights with external attention to fuse multi-view embeddings. Extensive experiments for two downstream tasks on real-world datasets demonstrate that our model outperforms state-of-the-art methods by up to 17% improvement .
@@
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; October 21–25, 2022; Birmingham, UKprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Neural networksccs: Computing methodologies Statistical relational learningccs: Information systems Data analyticsccs: General and reference Experimentation
@@ 1. Introduction
Urban region embedding is a classical embedding problem whose purpose is to to learn quantitative representations of regions from multi-sourced data. The problem has been found useful in many real-world applications such as socio-demographic feature prediction (Jean et al., 2018; Wang et al., 2020), crime prediction (Wang and Li, 2017; Zhang et al., 2020; Wu et al., 2022), economic growth prediction(Hui et al., 2020) and land usage classification (Yao et al., 2018; Zhang et al., 2020; Wu et al., 2022; Luo et al., 2022b).
@@
Urban region embedding is challenging due to the highly complex nature of urban data, deep learning approaches have attracted much interest in recent years. In particular, Graph neural networks (GNNs) have shown great potential for learning low-dimensional embeddings of graph-structured data (Cui et al., 2017; Abu-El-Haija et al., 2018; Kipf and Welling, 2016; Gilmer et al., 2017). Several existing studies that have attempted to integrate POI data and human mobility to characterize regions and obtain attractive results(Zhang et al., 2017; Wang and Li, 2017; Fu et al., 2019; Yao et al., 2018; Zhang et al., 2020; Chang et al., 2020; Liu et al., 2019; Wu et al., 2022; Luo et al., 2022a). For example, the methods (Zhang et al., 2020; Wu et al., 2022; Luo et al., 2022a) learns graph embeddings by combining multi-graph. The method(Zhang et al., 2020) employs an attention mechanism and simple crosstalk operations to aggregate information, while MFGN(Wu et al., 2022) deeply resolves the relevance of regions in fine-grained human mobility patterns. And the study(Luo et al., 2022a) aims to learn region representations that preserve both global-level inter-regional correlations, as well as local-level geographic context signals and inherent region properties. However, existing studies have predominantly focused on capturing correlations between neighboring regions, overlook the significant influence of distant regions in complex urban systems. Attention mechanism such as GAT(Veličković et al., 2018) assigns aggregation weights based solely on the influence of neighboring nodes, and self-attention(Vaswani et al., 2017) solely focuses on the regions themselves, both approaches fail to consider the potential correlations between regions and suffer from computational complexity. As a result, these methods yield sub-optimal embeddings and limit the model’s ability to capture underlying urban dynamics and features.
@@
To tackle these challenges, we propose a Region-Wise Multi-Graph Representation Learning (ROMER) for effective urban region representation with multi-view data. Our ROMER model adopts heterogeneous graph neural framework with respect to human movement patterns, POI semantics and check-in dynamics. In addition, a multi-graph aggregation module is design to capture region-wise dependencies and non-linear correlations among regions. Finally, we design an efficient and cost-effective dual-stage graph fusion module that learning adaptive weights for information sharing across diverse views with external attention and gating mechanism to fuse multi-view in an efficient and deeply collaborative manner.
@@ 2. Preliminaries
Definition 1. (Human mobility feature). The human mobility feature is defined as a trip sets $\mathcal{P}=\{p_{i}|p_{i}.{v_{o}},p_{i}.{v_{d}}\in V\}$ that occur in urban areas. $i=\{1,2,\cdots,M\}$, where $M$ is the number of trips.
Definition 2. (Semantic feature). The semantic feature describes the functional similarity among regions. Similar regions may necessarily be close in space. In this work, the semantic feature of the region is characterized with POIs in the located region. Given a region $v_{i}$, its semantic feature is represented as:
(1) $S=\left\{s_{i}\mid s_{i}\in\mathbb{R}\right\},i=1,2,\cdots,n$
@@
(2) $G=\left\{g_{i}\mid g_{i}\in\mathbb{R}\right\},i=1,2,\cdots,m$
Urban region embedding problem. The urban region embedding problem aims to learn a distributed and low dimensional embedding given the human mobility feature, semantic feature and dynamic feature. We denote the three features for a region $v_{i}$ as a vector $e^{i}\in\mathbb{R}^{r}$, where $r$ is the number of features. Then, the our final goal is to learn a mapping function $F$,
@@ 3. METHODOLOGY
The overall architecture of ROMER proposed in this paper is shown in Figure. 1, which consists of three components: the region-wise graph learning module, the multi-graph aggregation module and dual-stage fusion module.
@@ 3.1. Region Wise Graph Learning Module
3.1.1. Mobility based Region Graph
The movement of people within urban spaces across regions can be understood by examining their interactions. When people travel between different origins ($O$) and destinations ($D$), we can observe similarities in their patterns if they have the same O/D region. In simpler terms, by analysing the similarity of O/D patterns, we can identify important potential features related to human mobility. Given a set of human mobility $M$, we can use
@@ 3.1.1. Mobility based Region Graph
where $\mathcal{D}_{O}^{ij}$ denotes the dependencies between two origins, $\mathcal{D}_{D}^{ij}$ represents the dependencies between two destinations, $\operatorname{sim}(\cdot)$ denotes the cosine similarity. Based on Equation (6) and (7), we construct region wise graphs $\mathcal{G}_{O}$ and $\mathcal{G}_{D}$.
@@ 3.1.3. Dynamic Region Graph
where $\mathcal{D}_{G}^{ij}$ is the dynamic dependency between region $v_{i}$ and $v_{j}$. Now we construct dynamic region wise graph $\mathcal{G}_{G}$.
Figure 1. The architecture of ROMER consists of three key components: A) the region-wise graph learning module, B) the multi-graph aggregation module, and C) the dual-stage fusion module.
@@ 3.2. Multi-Graph Aggregation Module
It is observed that not only adjacent regions are relevant, but also many distant regions are correlated. However, existing GAT based methods (Zhang et al., 2020; Luo et al., 2022a) only considers the influence of neighbouring nodes. Inspired by (Huang et al., 2020) , we extract any two regions in the city that are correlated to assign learning weights. Given the vertex feature $\mathbf{h}=\left\{\vec{h}_{1},\vec{h}_{2},\ldots,\vec{h}_{n}\right\},\vec{h}_{i}\in\mathbb{R}^{F}$, where $F$ is the input dimension, the MGA layer works as follows:
@@
where $\sigma$ denotes the softmax function, $w_{ij}$ is the weight matrix. $A_{ij}$ is the similarity between $h_{i}$ and $h_{j}$. Softmax function is used to normalize the coefficients. $\lVert h_{i}\rVert$ denotes the norm of vector $h_{i}^{l}$ and $\cdot$ is the dot product of vectors.
@@
3.3. Dual-Stage Fusion Module
In this section, we design dual-stage fusion mechanism to efficient interchange information among multiple views, which works in two steps, attentive fusion followed by gated fusion.
3.3.1. Attentive Fusion
Self-attention is widely used in existing fusion methods(Vaswani et al., 2017), which is at the price of high computation. Inspired by (Guo et al., 2021), We adopt external attention to allow information to propagate across multiple views. Given the representations of $M$ views $\{\mathcal{E}_{1},\mathcal{E}_{2},\cdots,\mathcal{E}_{M}\}$, for each $\mathcal{E}_{i}$, we then propagate information among all views as follows:
@@ 3.3.1. Attentive Fusion
where $M_{k}\in\mathbb{R}^{S\times d}$ and $M_{v}\in\mathbb{R}^{S\times d}$ are learnable parameters independent of the representations, which act as the key and the value memory of the whole training dataset. And $d$ and $S$ are hyper-parameters. ^$\hat{\mathcal{E}}_{i}$ is considered as the relevant global information for $i$-th view. Meanwhile the normalization method we follow (Guo et al., 2021).
The embedding results ^$\hat{\mathcal{E}}_{O}$, ^$\hat{\mathcal{E}}_{D}$, ^$\hat{\mathcal{E}}_{S}$, and ^$\hat{\mathcal{E}}_{G}$ are generated from the above modules, linking the global information to subsequent gated fusions in the model.
3.3.2. Gated Fusion
This module is designed to integrate global and local area representation more effectively. To achieve this, We follow the fusion mechanism in (Zhang et al., 2020). The fusion layer operates as follows:
@@ 3.3.2. Gated Fusion
By incorporating the outputs of the base model into proposed joint learning module, we derive region embeddings ~$\tilde{\mathcal{E}}_{O},\tilde{\mathcal{E}}_{D},\tilde{\mathcal{E}}_{S}$, and ~$\tilde{\mathcal{E}}_{G}$, on which we work out of the various learning goals.
@@ 3.4. Prediction Objectives
Following (Zhang et al., 2020), given the source region $v_{i}$, we model the distribution of the target region $v_{j}$ as follows:
@@
Similarly, we model $\hat{p}_{D}\left(v_{j}\mid v_{i}\right)$ in the same way for the distribution of source region $r_{i}$ for a given destination region $r_{j}$. Then, $\mathcal{L}_{HMF}$ is constructed by maximizing the probability of O/D occurrence. Hence, the $\mathcal{L}_{HMF}$ between region $r_{i}$ and region $r_{j}$ can be computed as:
@@
In order for the learned region embeddings to preserve the region similarity across region attributes, we designed tasks to reconstruct region correlations based on the corresponding embeddings. Taking the Check-in property as an example, the learning objective is based on $\mathcal{C}_{G}$ and ~$\tilde{\mathcal{E}}_{G}=\left\{e_{G}^{i}\right\}_{i=1}^{n}$ defined as follows.
@@ 4.1. Datasets and Baselines
Experiments are conducted on several real-world datasets of New York City from NYC open data website [^1] https://opendata.cityofnewyork.us/. We follow (Zhang et al., 2020), taxi trip data was analysed as the source of human movement data, and the borough of Manhattan was divided into 180 zones to serve as the study area. As shown in Figure.  3(a), the borough of Manhattan was divided into 12 regions based on land use according to community board (Berg, 2007).
@@ 4.2. Experimental Settings
For land usage classification, we clustered the region embeddings using K-means with Normalized Mutual Information (NMI) and Adjusted Rand Index (ARI) with settings (Yao et al., 2018).In the case of check-in prediction, we utilize the Lasso regression (Tibshirani, 1996) with metrics of Mean Absolute Error (MAE), Root Mean Square Error (RMSE) and coefficient of determination ($R^{2}$).
@@ 4.3.2. Ablation Study
  • ROMER-G: It is ROMER without multi-graph aggregation Module(MGA). MGA module is replaced with the GAT(Veličković et al., 2018).
@@
  • ROMER-D: It is ROMER without Dual-stage graph fusion. The extracted spatial features are concatenated directly.
Figure. 2 displays the experimental results of ROMER and its variants in check-in prediction and land use classification tasks. The MGA module (ROMER-G) exhibits the most substantial impact on performance. Without the MGA component, the MAE for the prediction task increases significantly from 252.14 to 351.52, and the RMSE increases from 413.96 to 559.98. The attentive fusion module (ROMER-A) demonstrates the second-largest impact, affirming the effectiveness of our constructed module in enhancing single-view performance. Moreover, ROMER-D performs worse than ROMER, and the ablation study confirms the efficacy of the dual-stage fusion module in urban region embedding.
@@ 4.3.3. Visualized Analysis
To visually evaluate the clustering results, we plotted the clustering results of five methods in Figure. 3, where the same color marks the regions in the same cluster.We observe that the clustering results based on our method are optimal in terms of consistency with the real boundaries of ground conditions. These results suggest that the regional embeddings learned by our model are able to represent regional functions effectively.
Table 2. Computation time
Predicted Model Computation Time
Training (s/500epoch)
All downstream tasks MVGRE 166.45
MGFN 295.51
ROMER 142.22
4.4. Computation Time
In this section, on a workstation with Intel(R) Xeon(R) CPU E5-2686 v4 and NVIDIA GeForce RTX 3070 Ti-8G, we present the training time comparison of MVGRE(Zhang et al., 2020), MGFN(Wu et al., 2022), and ROMER in the tasks of check-in prediction and land usage classification. The training time results are summarized in Table 2. In particular, the results show that ROMER outperforms the other models, due to the replacement of the complex computational mechanism of self-attention(Vaswani et al., 2017).

### v2

@@ Abstract.
Urban region embedding is an important and yet highly challenging issue due to the complexity and constantly changing nature of urban data. To address the challenges, we propose a Region-Wise Multi-View Representation Learning (ROMER) to capture multi-view dependencies and learn expressive representations of urban regions without the constraints of rigid neighborhood region conditions. Our model focuses on learning urban region representation from multi-source urban data. First, we capture the multi-view correlations from mobility flow patterns, POI semantics, and check-in dynamics. Then, we adopt global graph attention networks to learn the similarity of any two vertices in graphs. To comprehensively consider and share features of multiple views, a two-stage fusion module is further proposed to learn weights with external attention to fuse multi-view embeddings. Extensive experiments for two downstream tasks on real-world datasets demonstrate that our model outperforms state-of-the-art methods by up to 17% improvement.
@@
journalyear: 2023copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Information and Knowledge Management; October 21–25, 2023; Birmingham, United Kingdombooktitle: Proceedings of the 32nd ACM International Conference on Information and Knowledge Management (CIKM ’23), October 21–25, 2023, Birmingham, United Kingdomprice: 15.00doi: 10.1145/3583780.3615194isbn: 979-8-4007-0124-5/23/10submissionid: shp2626ccs: Computing methodologies Neural networksccs: Information systems Clusteringccs: Information systems Data analytics
@@ 1. Introduction
Urban region embedding is a classical embedding problem whose purpose is to learn quantitative representations of regions from multi-sourced data. The problem has been found useful in many real-world applications such as socio-demographic feature prediction (Jean et al., 2018; Wang et al., 2020), crime prediction (Wang and Li, 2017; Zhang et al., 2020; Wu et al., 2022), economic growth prediction(Hui et al., 2020) and land usage classification (Yao et al., 2018; Zhang et al., 2020; Wu et al., 2022; Luo et al., 2022b).
@@
Urban region embedding is challenging due to the highly complex nature of urban data, deep learning approaches have attracted much interest in recent years. In particular, GNNs have shown great potential for learning low-dimensional embeddings of graph-structured data (Cui et al., 2017; Abu-El-Haija et al., 2018; Kipf and Welling, 2016; Gilmer et al., 2017; Wang et al., 2021b, a). Several existing studies that have attempted to integrate POI data and human mobility to characterize regions and obtain attractive results(Zhang et al., 2017; Wang and Li, 2017; Fu et al., 2019; Yao et al., 2018; Zhang et al., 2020; Chang et al., 2020; Liu et al., 2019; Wu et al., 2022; Luo et al., 2022a). For example, the methods (Zhang et al., 2020; Wu et al., 2022) learns graph embeddings by combining multi-graph. The method(Zhang et al., 2020) employs an attention mechanism and simple crosstalk operations to aggregate information, while MFGN(Wu et al., 2022) deeply resolves the relevance of regions in fine-grained human mobility patterns. However, existing studies have predominantly focused on capturing correlations between neighboring regions, overlooking the significant influence of distant regions in complex urban systems. Attention mechanism such as GAT(Veličković et al., 2018) assigns aggregation weights based solely on the influence of neighboring nodes, and self-attention(Vaswani et al., 2017) solely focuses on the regions themselves, both approaches fail to consider the potential correlations between regions and suffer from computational complexity(Guo et al., 2021). As a result, these methods produce sub-optimal embeddings and limit the ability of the model to capture the underlying urban dynamics and features.
@@
To tackle these challenges, we propose a Region-Wise Multi-Graph Representation Learning (ROMER) for effective urban region representation with multi-view data. Our ROMER model adopts a heterogeneous graph neural framework with respect to human movement patterns, POI semantics, and check-in dynamics. In addition, a multi-graph aggregation module is designed to capture region-wise dependencies and non-linear correlations among regions. Finally, we design an efficient and cost-effective attentive fusion module that learns adaptive weights for information sharing across diverse views with external attention and gating mechanisms to fuse multi-view in an efficient and deeply collaborative manner.
@@ 2. Preliminaries
Definition 1. (Human mobility feature). The human mobility feature is defined as a trip set $\mathcal{P}=\{p_{i}|p_{i}.{v_{o}},p_{i}.{v_{d}}\in V\}$ that occur in urban areas. $i=\{1,2,\cdots,M\}$, where $M$ is the number of trips.
Definition 2. (Semantic feature). The semantic feature describes the functional similarity among regions. Similar regions may necessarily be close in space. In this work, the semantic feature of the region is characterized by POIs in the located region. Given a region $v_{i}$, its semantic feature is represented as:
(1) $\mathcal{S}=\left\{s_{i}\mid s_{i}\in\mathbb{R}\right\},i=1,2,\cdots,n$
@@
(2) $\mathcal{G}=\left\{g_{i}\mid g_{i}\in\mathbb{R}\right\},i=1,2,\cdots,m$
Urban region embedding problem. We denote the three families of features for a region $v_{i}$ as a vector $e^{i}\in\mathbb{R}^{r}$, where $r$ is the number of features. Then, our final goal is to learn a mapping function $F$,
@@ 3. METHODOLOGY
The overall architecture of ROMER proposed in this paper is shown in Figure. 1, which consists of three components: the region-wise graph learning module, the multi-graph aggregation module, and the attentive fusion module.
@@ 3.1. Region Wise Graph Learning Module
3.1.1. Mobility-based Region Graph
The movement of people within urban spaces across regions can be understood by examining their interactions. When people travel between different origins ($O$) and destinations ($D$), we can observe similarities in their patterns if they have the same O/D region. In simpler terms, by analyzing the similarity of O/D patterns, we can identify important potential features related to human mobility. Given a set of human mobility $M$, we can use
@@ 3.1.1. Mobility-based Region Graph
where $\mathcal{D}_{O}^{ij}$ denotes the dependencies between two origins, $\mathcal{D}_{D}^{ij}$ represents the dependencies between two destinations, $\operatorname{sim}(\cdot)$ denotes the cosine similarity. Based on Equations (6) and (7), we construct region-wise graphs $\mathcal{G}_{O}$ and $\mathcal{G}_{D}$.
@@ 3.1.3. Dynamic Region Graph
where $\mathcal{D}_{G}^{ij}$ is the dynamic dependency between region $v_{i}$ and $v_{j}$. Now we construct dynamic region-wise graph $\mathcal{G}_{G}$.
Figure 1. The architecture of ROMER consists of three key components: A) the region-wise graph learning module, B) the multi-graph aggregation module, and C) the attentive fusion module.
@@ 3.2. Multi-Graph Aggregation Module
It is observed that not only adjacent regions are relevant, but also many distant regions are correlated. However, existing GAT-based methods (Zhang et al., 2020; Luo et al., 2022a) only consider the influence of neighboring nodes. Inspired by (Huang et al., 2020), we utilize an improved GAT mechanism to extract any two relevant regions in a city to assign learning weights. Given the vertex feature $\mathbf{h}=\left\{\vec{h}_{1},\vec{h}_{2},\ldots,\vec{h}_{n}\right\},\vec{h}_{i}\in\mathbb{R}^{F}$, where $F$ is the input dimension, the MGA layer works as follows:
@@
where $\sigma$ denotes the softmax function, $w_{ij}$ is the weight matrix. $A_{ij}$ is the similarity between $h_{i}$ and $h_{j}$. The softmax function is used to normalize the coefficients. $\lVert h_{i}\rVert$ denotes the norm of vector $h_{i}^{l}$ and $\cdot$ is the dot product of vectors.
@@
3.3. Attentive Fusion Module
In this section, we design an attentive fusion mechanism to efficiently interchange information among multiple views. The self-attention mechanism is widely used in existing fusion methods(Vaswani et al., 2017), which is at the price of high computation. Inspired by (Guo et al., 2021), We adopt external attention to allow information to propagate across multiple views. Given the representations of $M$ views $\{\mathcal{E}_{1},\mathcal{E}_{2},\cdots,\mathcal{E}_{M}\}$, for each $\mathcal{E}_{i}$, we then propagate information among all views as follows:
@@ 3.3. Attentive Fusion Module
where $M_{k}\in\mathbb{R}^{S\times d}$ and $M_{v}\in\mathbb{R}^{S\times d}$ are learnable parameters independent of the representations, which act as the key and the value memory of the whole training dataset. And $d$ and $S$ are hyper-parameters. ^$\hat{\mathcal{E}}_{i}$ is considered as the relevant global information for $i$-th view. Meanwhile, the normalization method we follow (Guo et al., 2021). The embedding results ^$\hat{\mathcal{E}}_{O}$, ^$\hat{\mathcal{E}}_{D}$, ^$\hat{\mathcal{E}}_{S}$, and ^$\hat{\mathcal{E}}_{G}$ are generated from the above modules, linking the global information to subsequent fusions in the model.
To integrate global and local region representation. We follow the fusion mechanism in (Zhang et al., 2020). The fusion layer operates as follows:
@@
By incorporating the outputs of the base model into the proposed joint learning module, we derive region embeddings ~$\tilde{\mathcal{E}}_{O},\tilde{\mathcal{E}}_{D},\tilde{\mathcal{E}}_{S}$, and ~$\tilde{\mathcal{E}}_{G}$, on which we work out the various learning goals.
@@ 3.4. Prediction Objectives
In order to efficiently train our model, according to (Zhang et al., 2020), given the source region $v_{i}$, we model the distribution of the target region $v_{j}$ as follows:
@@
Similarly, we model $\hat{p}_{D}\left(v_{j}\mid v_{i}\right)$ similarly for the distribution of source region $r_{i}$ for a given destination region $r_{j}$. Then, $\mathcal{L}_{HMF}$ is constructed by maximizing the probability of O/D occurrence. Hence, the $\mathcal{L}_{HMF}$ between region $r_{i}$ and region $r_{j}$ can be computed as:
@@
For the learned region embeddings to preserve the region similarity across region attributes, we designed tasks to reconstruct region correlations based on the corresponding embeddings. Taking the Check-in property as an example, the learning objective is based on $\mathcal{C}_{G}$ and ~$\tilde{\mathcal{E}}_{G}=\left\{e_{G}^{i}\right\}_{i=1}^{n}$ defined as follows.
@@ 4.1. Datasets and Baselines
Experiments are conducted on several real-world datasets of New York City from NYC open data website [^1] https://opendata.cityofnewyork.us/. We follow (Zhang et al., 2020), taxi trip data was analyzed as the source of human movement data, and the borough of Manhattan was divided into 180 zones to serve as the study area. As shown in Figure.  3(a), the borough of Manhattan was divided into 12 regions based on land use according to community board (Berg, 2007).
@@ 4.2. Experimental Settings
For land usage classification, we clustered the region embeddings using K-means with Normalized Mutual Information (NMI) and Adjusted Rand Index (ARI) with settings (Yao et al., 2018).In the case of check-in prediction, we utilize the Lasso regression (Tibshirani, 1996) with metrics of Mean Absolute Error (MAE), Root Mean Square Error (RMSE), and coefficient of determination ($R^{2}$).
@@ 4.3.2. Ablation Study
  • ROMER-G: It is ROMER without a multi-graph aggregation Module (MGA). MGA is replaced with the GAT(Veličković et al., 2018).
@@
  • ROMER-D: It is ROMER without Attentive graph fusion. The extracted spatial features are concatenated directly.
Figure. 2 displays the experimental results of ROMER and its variants in check-in prediction and land use classification tasks. The MGA module (ROMER-G) exhibits the most substantial impact on performance. Without the MGA component, the MAE for the prediction task increases significantly from 252.14 to 351.52, and the RMSE increases from 413.96 to 559.98. The attentive fusion module (ROMER-A) demonstrates the second-largest impact, affirming the effectiveness of our constructed module in enhancing single-view performance.
@@ 4.3.3. Visualized Analysis
To visually evaluate the clustering results, we plotted the clustering results of five methods in Figure. 3, where the same color marks the regions in the same cluster. We observe that the clustering results based on our method are optimal in terms of consistency with the real boundaries of ground conditions. These results suggest that the regional embeddings learned by our model can represent regional functions effectively.